News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 3 October 2003 7.30am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was completed on Oct 1 to uplink a flight software k-constant patch to update the IRU-2 calibration constants. The update is based on data taken since the IRU swap in July and initial data following the uplink indicates the expected improvement in maneuver 1-shot updates.

An anomaly related to the HRC was detected on Sep 26. During the commanding prior to the start of an observation the HRC -Y shutter failed to insert. Review of the telemetry indicates that the HRC did not select the -Y shutter motor when commanded for either the insert or the subsequent retraction. The anomaly is not expected to impact other observations in the short term since the shutters are rarely used. An anomaly team has been formed and is performing fault tree analysis, contingency planning and motor control testing on the proof of concept model.

A Chandra press release was issued on Oct 1 describing observations of a giant elliptical galaxy moving at high speed through the diffuse hot gas within the Virgo galaxy cluster. For details see: http://chandra.harvard.edu/photo/2003/m86/.

The observing schedule for next week is shown below and includes an ACA dark current calibration.

------------------------------------------
4C 74.26               ACIS-S/HETG  Oct 6
Radiation Belts
MACS-J0647.7+7015      ACIS-I       Oct 7
B1600+434              ACIS-S
4C 74.26               ACIS-S/HETG
1735-269               ACIS-S       Oct 8
PSR B1046-58           ACIS-S
Radiation Belts                     Oct 9
4U1254-690             ACIS-S/HETG
Abell 2244             ACIS-S       Oct 10
G70.7+1.2              ACIS-S       Oct 11
A4038                  ACIS-I
ACA Dark Current Cal

------------------------------------------

All spacecraft subsystems with the exception of the HRC as noted above continued to operate nominally.

Return to Status Reports